OpenStudy (anonymous):

Is sin of theta = 2/3 with theta in quadrant 1, what is sec of theta?

OpenStudy (anonymous):

use this to find cos\(\theta\): \(sin^2\theta+cos^2\theta=1\) later: sec\(\theta\)=1/cos\(\theta\)